OpenStudy (anonymous):

|x|+5=8 SOLVE A {-13, -3} B {-13} C {-3, 3} D {13, -3}

OpenStudy (anonymous):

@texaschic101 @d3Xter

OpenStudy (anonymous):

I think it is C

OpenStudy (anonymous):

Wrong

OpenStudy (anonymous):

then is it D or B

OpenStudy (anonymous):

Are you just guessing? This seems more or less like an inequality problem...

OpenStudy (anonymous):

no I just suck at these

OpenStudy (anonymous):

Are you just guessing? This seems more or less like an inequality problem...\[\left| x + 5 \right| \le 8 \rightarrow -8 \le | x+5|≤8 \rightarrow x \le (-13,3)\]

OpenStudy (anonymous):

OHHHH told you I sucked at this